What quantity of 73% chlorine solution must be mixed with a 30% chlorine solution to produce 399.12 mL of 69% solution?

Answer by Alan3354(69443)

What quantity of 73% chlorine solution must be mixed with a 30% chlorine solution to produce 399.12 mL of 69% solution?
---------
s + t = 399.12 (total solution)
73s + 30t = 399.12*69